# From small eigenvalues to large cuts, and Chowla’s cosine problem

### v1

@@ Abstract
We show that there exists an absolute constant $\gamma>0$ such that for every $A\subseteq\mathbb{Z}_{>0}$ we have
$\min_{x\in[0,2\pi]}\sum_{a\in A}\cos(ax)\leq-\Omega(|A|^{\gamma}).$
This gives the first polynomial bound for Chowla’s cosine problem from 1965. To show this, we prove structural statements about graphs whose smallest eigenvalue is small in absolute value. As another application, we show that any graph $G$ with $m$ edges and no clique of size $m^{1/2-\delta}$ has a cut of size at least $m/2+m^{1/2+\varepsilon}$ for some $\varepsilon=\varepsilon(\delta)>0$. This proves a weak version of a celebrated conjecture of Alon, Bollobás, Krivelevich, and Sudakov. Our proofs are based on novel spectral and linear algebraic techniques, involving subspace compressions and Hadamard products of matrices.
@@ 1 Introduction
The central question in spectral graph theory is how the structural properties of a graph influence its spectrum and, conversely, what information about the graph the spectrum encodes. Here and throughout, the spectrum of the graph refers to the set of eigenvalues of its adjacency matrix. Some of the most prominent relations between eigenvalues and structural properties of a graph are the Expander Mixing Lemma [4] (which connects pseudorandomness properties of the graph to its eigenvalues) and the Hoffman bound [43] (which bounds the size of the largest independent set based on the smallest eigenvalue). The question which concerns us in this paper is the following: what can be said about the graphs whose last eigenvalue is small in absolute value?
The prime examples of such graphs are cliques, since all of their eigenvalues are at least 1-1. As taking disjoint unions of graphs corresponds to taking a union of their spectra, and changing a small number of edges does not affect the spectrum significantly, it is not hard to see that graphs which are close to disjoint unions of cliques also have small $|\lambda_{n}|$, where $\lambda_{n}$ denotes the smallest eigenvalue of the graph. The main technical result of our paper is a converse to this statement: we prove that already a mild restriction $|\lambda_{n}|\leq n^{1/4-o(1)}$ on the smallest eigenvalue $\lambda_{n}$ of an $n$-vertex graph forces the graph to be close to a disjoint union of cliques, and the exponent $1/4$ is best possible.
We also obtain a variant of this result which applies to sparse graph. More precisely, we show that graphs of average degree $d$ and smallest eigenvalue $|\lambda_{n}|\leq d^{\gamma}$ contain cliques of size $d^{1-O(\gamma)}$. Perhaps surprisingly, this simple graph-theoretic statement has a powerful application to Chowla’s cosine problem [21] from 1965. In fact, it allows us to prove that if $A$ is a finite set of positive integers, then the function
@@
takes values as small as $|A|^{\Omega(1)}$. For a more detailed overview of Chowla’s cosine problem and related work, consult Section 1.1.
Our methods can also be used to study graphs with small maximum cut, allowing us to make substantial progress on a celebrated conjecture of Alon, Bollobás, Krivelevich, and Sudakov [3]. This is a central problem in the study of the maximum cut in $H$-free graphs, initiated by Erdős and Lovász [35] in the 1970’s. If $H$ is a fixed graph, this conjecture states that any graph $G$ with $m$ edges and no subgraph isomorphic to $H$ has a cut of size at least $m/2+m^{1/2+\varepsilon_{H}}$ for some $\varepsilon_{H}>1/4$. Here, a cut is a partition of a vertex set of $G$ into two parts, and its size is the number of edges crossing the partition. We prove that forbidding a clique of size as large as $m^{1/2-\delta}$ in $G$ already allows to show that $G$ has a cut with at least $m/2+m^{1/2+\varepsilon}$ edges, for some $\varepsilon=\varepsilon(\delta)>0$. The problem of estimating the size of the maximum cut is discussed in further detail in Section 1.2.
Finally, our results have further implications about graphs of small second eigenvalue. A classical result of Alon and Boppana [63] gives a bound on the second-largest eigenvalue $\lambda_{2}$ of $d$-regular $n$-vertex graphs, and it states that $\lambda_{2}\geq 2\sqrt{d-1}(1-\frac{1}{\lfloor D/2\rfloor})$, where $D$ is the diameter of the graph. Note that this bound is trivial for graphs of diameter at most $3$, which could happen already when $d\geq n^{1/3}$. In this paper, we extend Alon–Boppana theorem to dense graphs, and we show that if an $n$-vertex regular graph is far from a Turán graph, then its second eigenvalue satisfies $\lambda_{2}\geq n^{1/4-o(1)}$, with the exponent $1/4$ being optimal. Finally, we prove similar structural results about graphs of large bisection width. Our proofs introduce novel spectral and linear-algebraic techniques based on subspace compressions of matrices and the use of Hadamard products, which may be of independent interest.
@@ 1.1 Chowla’s cosine problem
In 1948, in the study of certain Dedekind zeta functions, Ankeny and Chowla came across the the following question, see [20] — is it true that for every $K>0$ and sufficiently large $n>0$, if $a_{1},\dots,a_{n}$ are distinct integers, then the minimum of the function $f(x)=\cos(a_{1}x)+\cdots+\cos(a_{n}x)$ is less than K-K ? This was proved by Uchiyama and Uchiyama [74], albeit with poor quantitative dependencies. Soon after this work, in 1965, Chowla [21] revisited the problem and made a more precise conjecture, today known as Chowla’s cosine problem. He asked to show that $\min_{x\in[0,2\pi]}f(x)=\min_{x\in[0,2\pi]}\sum_{i=1}^{n}\cos(a_{i}x)\leq-\Omega(\sqrt{n})$. The bound $\Omega(\sqrt{n})$ comes from the fact that in case $A=\{a_{1},\dots,a_{n}\}$ can be written as A=BBA=B-B, where $B$ is a Sidon set, one has $\min_{x\in[0,2\pi]}f(x)=-\Theta(\sqrt{n})$, see [62] for a detailed proof. The bounds of Uchiyama and Uchiyama were later improved by Roth [67], who showed a lower bound of $\min_{x}f(x)\leq-\Omega(\sqrt{\log n/\log\log n})$.
An important observation in the early study of Chowla’s cosine problem was its connection to Littlewood’s $L_{1}$-problem, which asks to show that for each $n$-element set $A\subseteq\mathbb{Z}$, the $L_{1}$-norm of the Fourier transform of $\mathds{1}_{A}$ is bounded below by $\log n$, i.e.
@@
Indeed, any lower bound on Littlewood’s $L_{1}$-problem gives a comparable upper bound for the cosine problem (see [67] for a detailed derivation). Thus, the resolution of the Littlewood $L_{1}$-problem by Konyagin [53] and McGehee, Pigno and Smith [61] immediately implies that $\min_{x}f(x)\leq-\Omega(\log n)$.
It was Bourgain [14] who first broke this logarithmic barrier, and then his method was further refined by Ruzsa [68] to give the previously best known bound $\min_{x}f(x)=-\exp\big{(}\Omega(\sqrt{\log n})\big{)}$. Chowla’s cosine problem also appears on Green’s 100 problems list [41] as problem number 81. Here, we give the first polynomial bound.
@@
There exists an absolute constant $\gamma>0$ such that for any $A$ of positive integers, there exists $x\in[0,2\pi]$ such that
$\sum_{a\in A}\cos(ax)\leq-\Omega(|A|^{\gamma}).$
Following our proof, one can take $\gamma=0.01$, so it remains an interesting open problem to decide whether $\gamma=1/2$ is the best exponent. Our graph theoretic machinery has a hard theoretical barrier at the exponent $1/4$, and possibly substantial new ideas are needed to move beyond this point (however, we do not claim that $\gamma=1/4$ is reachable either with our methods). We now say a couple of words about the proof of Theorem 1.1. The key ingredient of the proof is the following purely graph-theoretic result.
@@
For every $\gamma>0$, the following holds for every sufficiently large $d$. Let $G$ be a graph of average degree $d$ and assume that $|\lambda_{n}|\leq d^{\gamma}$. Then $G$ contains a clique of size at least $d^{1-O(\gamma)}$.
The first idea is to embed $A$ into the group $\mathbb{Z}/n\mathbb{Z}$ for a sufficiently large prime $n$ and to consider the Cayley graph $\Gamma=\mbox{Cay}(\mathbb{Z}/n\mathbb{Z},A\cup-A)$. It is well-known that the eigenvalues of Cayley graphs correspond to the Fourier coefficients of the generating set, and thus the smallest eigenvalue $\lambda_{n}$ satisfies $\lambda_{n}=\sum_{a\in A\cup-A}\exp(2\pi iak/n)=2\sum_{a\in A}\cos(2\pi ak/n)$ for some $k\in[n]$. Hence, $\frac{1}{2}\lambda_{n}\geq\min_{x}f(x)$. In other words, unless the minimum of $f(x)$ is very negative, $|\lambda_{n}|$ is small. But then we can use Theorem 1.2 to conclude that $G$ contains very large cliques. However, the transitive symmetry of the Cayley graph does not allow such large cliques, without violating the property of having small $|\lambda_{n}|$.
We conclude the discussion of Chowla’s cosine problem by remarking some other interesting questions about cosine polynomials. For example, the problem of estimating the number of zeros of such polynomials have also attracted lots of attention recently. Namely, an old problem of Littlewood [59] asks study the minimum number of zeros of the function $f(x)=\sum_{i=1}^{n}\cos(a_{i}x)$ in the interval $0,2\pi]$. Although Littlewood guessed that this number should always be linear in $n$, Borwein, Erdélyi, Ferguson and Lockhart [15] showed that the there are integers $a_{1},\dots,a_{n}$ such that $f(x)$ has at most $n^{5/6}$ zeros. This result was later improved to $(n\log n)^{2/3}$ by Juškevičius and Sahasrabudhe [49] and, independently, by Konyagin [54]. A complementary bound has been proven by Sahasrabudhe [69] and Erdélyi [33, 34], who showed that $f(x)$ always has at least $(\log\log\log n)^{1/2-\varepsilon}$ roots, and this was later improved to $(\log\log n)^{1-o(1)}$ by Bedert [13].
In the Appendix, we discuss further extensions of Chowla’s problem in arbitrary finite groups. The proof of Theorem 1.1 is presented in Section 3, and the proof of Theorem 1.2 is presented in Section 12.
@@ 1.2 Maximum Cut
Given a graph $G$, a cut in $G$ is a partition $(U,V)$ of the vertex set together with all the edges having exactly one endpoint in both parts. The size of the cut is the number of its edges. The MaxCut of $G$ is the maximum size of a cut, denoted by $\mbox{mc}(G)$. The MaxCut is among one the most studied graph parameters, lying at the intersection of theoretical computer science [32, 40, 51], extremal combinatorics [2, 3, 30, 35] and probabilistic graph theory [23, 22, 27]. In theoretical computer science, one is usually interested in approximating the size of the MaxCut efficiently, and in extremal combinatorics the goal is often to establish good bounds on the MaxCut depending on various graph parameters, such as the number of edges or vertices of the graph.
@@
A simple probabilistic argument shows that every graph with $m$ edges has a cut of size at least $m/2$. Indeed, a random cut, chosen from the uniform distribution on all cuts, has size $m/2$ in expectation. The constant $1/2$ cannot be improved in general, and therefore when measuring the size of the MaxCut in $G$, it is often more natural to talk about the surplus of $G$, which is defined as $\operatorname{surp}(G)=\mbox{mc}(G)-m/2$. The trivial bound can be improved, a fundamental result of Edwards [30, 31] shows that any graph $G$ with $m$ edges has $\mbox{mc}(G)\geq\frac{m}{2}+\frac{\sqrt{8m+1}-1}{8}$ or, equivalently, that $\operatorname{surp}(G)\geq\frac{\sqrt{8m+1}-1}{8}$, which is sharp when $G$ is a clique on an odd number of vertices.
In general, if $G$ is a disjoint union of constantly many cliques, then the MaxCut of $G$ is of size $m/2+O(\sqrt{m})$. This raises the following natural question.
Can this bound be improved if $G$ is far from a disjoint union of cliques?
@@
One way to ensure that a graph is far from a disjoint union of cliques is to assume that it does not contain some fixed graph $H$ as a subgraph. The study of the size of the MaxCut, and in turn the surplus, in such graphs was initiated by Erdős and Lovász in the 1970’s (see [35]). One of the first major results in the area is due to Alon [2], who proved that if a graph $G$ has $m$ edges and no triangles, then $\operatorname{surp}(G)=\Omega(m^{4/5})$, and this bound is tight. There are two natural ways to generalize this result - one is to study graphs without short cycles and the other is to study graphs avoiding $K_{r}$, the complete graph on $r$ vertices.
The surplus in graphs without short cycles, studied in [3, 6, 10, 39], proved to be the easier of these two problems, and [10, 39] achieve tight bounds for this problem. On the other hand, finding the size of the minimum surplus in $K_{r}$-free graphs seems to be much more difficult. Alon, Bollobás, Krivelevich, and Sudakov [6] proved that for every $r$, there exists $\varepsilon_{r}>0$ such that every $K_{r}$-free graph has surplus at least $m^{1/2+\varepsilon_{r}}$. This was improved by Carlson, Kolla, Li, Mani, Sudakov, and Trevisan [18], and then Glock, Janzer, and Sudakov [39] established $\operatorname{surp}(G)\geq m^{\frac{1}{2}+\frac{3}{4r-2}}$.
@@
However, these results seem to be far from the truth, since Alon, Bollobás, Krivelevich and Sudakov conjectured in [3] that the answer should be $\operatorname{surp}(G)\geq m^{3/4+\varepsilon_{r}}$ for some $\varepsilon_{r}>0$. This conjecture is still wide open, and, in fact, for a long time, it was a tantalizing open problem to find any absolute constant $\varepsilon>0$ (independent of $r$), such that every $K_{r}$-free graph has surplus $\Omega_{r}(m^{1/2+\varepsilon})$. Glock, Janzer and Sudakov [39] write “Arguably, the main open problem is to decide whether there exists a positive absolute constant $\varepsilon$ such that any $K_{r}$-free graph with $m$ edges has surplus $\Omega_{r}(m^{1/2+\varepsilon})$.” Our next main result not only proves this, but shows that we can achieve such a large surplus by forbidding extremely large cliques as well.
@@
Coming back to the motivating question of whether graphs with small surplus must necessarily look like unions of cliques, we also prove the following stability result.

### v2

@@ Abstract
We prove that every graph with average degree $d$ and smallest adjacency eigenvalue $|\lambda_{n}|\leq d^{\gamma}$ contains a clique of size $d^{1-O(\gamma)}$. A simple corollary of this yields the first polynomial bound for Chowla’s cosine problem (1965): for every finite set $A\subseteq\mathbb{Z}_{>0}$, the minimum of the cosine polynomial satisfies
$\min_{x\in[0,2\pi]}\sum_{a\in A}\cos(ax)\leq-|A|^{1/10-o(1)}.$
Another application makes significant progress on the problem of MaxCut in $H$-free graphs initiated by Erdős and Lovász in the 1970’s. We show that every $m$-edge graph with no clique of size $m^{1/2-\delta}$ has a cut of size at least $m/2+m^{1/2+\varepsilon}$ for some $\varepsilon=\varepsilon(\delta)>0$.
@@ 1 Introduction
A central theme in spectral graph theory is understanding the interplay between the structural properties of a graph and its spectrum. Here and throughout, the spectrum of the graph refers to the set of eigenvalues of its adjacency matrix. Some of the most prominent results highlighting the connection between eigenvalues and structural properties are the expander mixing lemma [4], which relates eigenvalues to pseudorandomness, and Hoffman’s bound [50], which uses the smallest eigenvalue to bound the independence number.
In this paper, we study graphs whose smallest eigenvalue is not very negative. The prime examples of such graphs are complete graphs, where every eigenvalue is at least 1-1. In fact, all graphs that are close to disjoint unions of cliques have their smallest eigenvalue small in absolute value. Our main technical contribution is a converse to this observation: we prove that even the mild restriction $|\lambda_{n}|\leq n^{1/4-o(1)}$ on the smallest eigenvalue $\lambda_{n}$ of an $n$-vertex graph forces the graph to be close to a disjoint union of cliques. Moreover, the exponent $1/4$ is best possible.
We further obtain an analogue of this result for sparse graphs. We show that graphs of average degree $d$ and smallest eigenvalue $|\lambda_{n}|\leq d^{\gamma}$ contain cliques of size $d^{1-O(\gamma)}$. Perhaps surprisingly, this purely graph-theoretic statement has a powerful application to Chowla’s cosine problem [24], a classical problem from harmonic analysis. It implies that if $A$ is a finite set of positive integers, then the cosine polynomial
@@
attains values as small as $|A|^{\Omega(1)}$. A more detailed discussion of Chowla’s cosine problem is presented in Section 1.1.
Our methods are also applicable in the study of graphs with small maximum cut, leading to substantial progress on a celebrated conjecture of Alon, Bollobás, Krivelevich, and Sudakov [3]. This is a central problem in the study of the maximum cut in $H$-free graphs, a topic initiated by Erdős and Lovász [40] in the 1970’s. The conjecture asserts that for any fixed graph $H$, every $H$-free graph $G$ with $m$ edges has a cut of size at least $m/2+m^{3/4+\varepsilon_{H}}$ for some $\varepsilon_{H}>0$. Over the decades, even the weaker question of whether one can guarantee a cut of size at least $m/2+m^{1/2+\varepsilon}$, for some fixed $\varepsilon>0$, remained wide open. We not only resolve this, but also prove a stronger statement: if $G$ contains no clique of size $m^{1/2-\delta}$, then $G$ admits a cut of size at least
$m/2+m^{1/2+\varepsilon}$
for some $\varepsilon=\varepsilon(\delta)>0$. A more detailed discussion of the maximum cut problem is provided in Section 1.2.
Finally, our results have further implications about the second eigenvalue. A classical theorem of Alon and Boppana [73] lower bounds the second-largest eigenvalue $\lambda_{2}$ of $d$-regular $n$-vertex graphs by
$\lambda_{2}\geq 2\sqrt{d-1}\left(1-\frac{1}{\lfloor D/2\rfloor}\right),$
where $D$ is the diameter of the graph. This estimate becomes trivial for $D\leq 3$, which can already occur for $d\gg n^{1/3}$. We extend the Alon-Boppana bound to dense graphs: we show that if a regular graph is far from being a Turán graph, then
$\lambda_{2}\geq n^{1/4-o(1)}$
and the exponent $1/4$ is best possible.
Our proofs introduce novel spectral and linear-algebraic techniques based on subspace compressions of matrices and the use of Hadamard products, which may be of independent interest.
@@ 1.1 Chowla’s cosine problem
In 1948, in the study of certain Dedekind zeta functions, Ankeny and Chowla came across the following question (see [23]): is it true that for every $K>0$ and sufficiently large $n>0$, if $a_{1},\dots,a_{n}$ are distinct positive integers, then the minimum of the function $f(x)=\cos(a_{1}x)+\cdots+\cos(a_{n}x)$ is less than K-K? Soon thereafter, Uchiyama and Uchiyama [85] answered this question affirmatively, but with poor quantitative dependencies, by observing the connection to Cohen’s work [25] on Littlewood’s $L_{1}$-problem. This problem asks to show that for each $n$-element set $A\subseteq\mathbb{Z}$, the $L_{1}$-norm of the Fourier transform of $\mathds{1}_{A}$ is bounded below by $\Omega(\log n)$, i.e.
@@
Any lower bound on Littlewood’s $L_{1}$-problem gives a comparable upper bound for Chowla’s cosine problem, see [77] for a detailed derivation.
In 1965, Chowla [24] revisited the problem and made a more precise conjecture, today known as Chowla’s cosine problem: show that for an $n$-element set $A$ of positive integers,
$\min_{x\in[0,2\pi]}f(x)=\min_{x\in[0,2\pi]}\sum_{a\in A}\cos(ax)\leq-\Omega(\sqrt{n}).$
In case $A$ can be written as A=BBA=B-B, where $B$ is a Sidon set, one has $\min_{x\in[0,2\pi]}f(x)=-\Theta(\sqrt{n})$ (see [71] for a detailed proof), so if the conjecture is true, the bound $\Omega(\sqrt{n})$ is the best possible.
The subsequent decades saw a persistent interest in this problem, and the bounds of Uchiyama and Uchiyama were improved by Roth [77] in 1973, who showed that $\min_{x}f(x)\leq-\Omega(\sqrt{\log n/\log\log n})$. Then the resolution of the Littlewood $L_{1}$-problem in 1981 by Konyagin [61] and McGehee, Pigno and Smith [70] improved this to $\min_{x}f(x)\leq-\Omega(\log n)$. It was Bourgain [17] who first broke this logarithmic barrier, and his method was further refined in 2004 by Ruzsa [78] to give the previously best known bound $\min_{x}f(x)\leq-\exp\big(\Omega(\sqrt{\log n})\big)$. Chowla’s cosine problem is also highlighted as problem number 81 on Green’s 100 problems list [48]. Here, we give the first polynomial bound.
@@
For any finite set $A$ of positive integers, there exists $x\in[0,2\pi]$ such that
$\sum_{a\in A}\cos(ax)\leq-|A|^{1/10-o(1)}.$
We now say a few words about the proof of Theorem 1.1. The key ingredient of the proof is the following graph-theoretic result.
@@
For every $\gamma\in(0,1/10)$, the following holds for every sufficiently large $d$. Let $G$ be a graph of average degree $d$ and assume that the smallest eigenvalue $\lambda_{n}$ of $G$ satisfies $|\lambda_{n}|\leq d^{\gamma}$. Then $G$ contains a clique of size at least $d^{1-4\gamma}$.
We embed $A$ into the group $\mathbb{Z}/n\mathbb{Z}$ for a sufficiently large prime $n$ and consider the Cayley graph $G=\operatorname{Cay}(\mathbb{Z}/n\mathbb{Z},A\cup-A)$. It is well-known that the eigenvalues of Cayley graphs correspond to the Fourier coefficients of the generating set, and thus the smallest eigenvalue $\lambda_{n}$ satisfies
$\lambda_{n}=\sum_{a\in A\cup-A}\exp(2\pi iak/n)=2\sum_{a\in A}\cos(2\pi ak/n)$
for some $k\in\mathbb{Z}/n\mathbb{Z}$. Hence, $\frac{1}{2}\lambda_{n}\geq\min_{x}f(x)$, so a lower bound on $\min_{x}f(x)$ yields an upper bound on $|\lambda_{n}|$. Then by Theorem 1.2, $G$ contains very large cliques. However, by appealing to the transitive symmetry of the Cayley graph, we show that the existence of large cliques forces large $|\lambda_{n}|$.
We conclude this section by noting that cosine polynomials are the subject of a number of other interesting problems. An old problem of Littlewood [68] asks to study the minimum number of zeros of $f(x)=\sum_{a\in A}\cos(ax)$ in the interval $0,2\pi]$, where $A$ is a set of $n$ positive integers. Although Littlewood conjectured that this number is linear in $n$, Borwein, Erdélyi, Ferguson and Lockhart [18] showed that there are integers $a_{1},\dots,a_{n}$ such that $f(x)$ has at most $n^{5/6+o(1)}$ zeros. This result was later improved to $O((n\log n)^{2/3})$ by Juškevičius and Sahasrabudhe [57] and, independently, by Konyagin [62]. A complementary bound has been proven by Sahasrabudhe [79] and Erdélyi [37, 38], who showed that $f(x)$ always has at least $(\log\log\log n)^{1/2-\varepsilon}$ roots, which was later improved to $(\log\log n)^{1-o(1)}$ by Bedert [15]. Another problem about trigonometric polynomials, asked by Littlewood [67] and Erdős [39], concerns the existence of “flat” polynomials, i.e. polynomials $f(z)=\sum_{k=0}^{n}\varepsilon_{k}z^{k}$ with coefficients $\varepsilon_{k}=\pm 1$ such that $|f(z)|=\Theta(\sqrt{n})$ for all $|z|=1$. Writing $z=\cos\theta+i\sin\theta$ shows that this problem is about controlling the size of trigonometric polynomials $f(z)=\sum_{k=0}^{n}\varepsilon_{k}(\cos k\theta+i\sin k\theta)$. The existence of such functions was proved only very recently by Balister, Bollobás, Morris, Sahasrabudhe, and Tiba [9].
In Appendix˜A, we discuss further extensions of Chowla’s problem in arbitrary finite groups. The proof of Theorem 1.1 is presented in Section 3, and the proof of Theorem 1.2 is presented in Section 10.
Note added after publication. Very recently, Bedert [16] achieved a result similar to our Theorem˜1.1, proving that for every $A\subset\mathbb{Z}_{>0}$, one has $\min_{x}\sum_{a\in A}\cos(ax)\leq-\Omega(|A|^{1/7-o(1)})$. Interestingly, the methods of Bedert are fundamentally different from ours. The proof in [16] is Fourier-analytic, compared to our spectral and graph-theoretic approach.
@@ 1.2 Maximum Cut
Given a graph $G$, a cut in $G$ is a partition $(U,V)$ of the vertex set together with all the edges having exactly one endpoint in both parts. The size of the cut is the number of its edges. The maximum cut (or MaxCut) of $G$, denoted by $\mbox{mc}(G)$, is the largest possible size of a cut. The MaxCut is among the most extensively studied graph parameters, lying at the intersection of theoretical computer science [36, 47, 59], extremal combinatorics [2, 3, 34, 40] and probabilistic graph theory [27, 26, 31]. In theoretical computer science, one is usually interested in efficient approximations of the maximum cut, while in extremal combinatorics the emphasis is on establishing sharp bounds in terms of various graph parameters, such as the number of vertices or edges.
@@
A simple probabilistic argument shows that every graph with $m$ edges has a cut of size at least $m/2$. Indeed, if one chooses a cut uniformly at random, its expected size is exactly $m/2$. The constant $1/2$ is best possible in general, and therefore it is often more natural to measure the surplus of a graph $G$, defined as $\operatorname{surp}(G)=\mbox{mc}(G)-m/2$. A fundamental result of Edwards [34, 35] asserts that any graph $G$ with $m$ edges satisfies $\mbox{mc}(G)\geq\frac{m}{2}+\frac{\sqrt{8m+1}-1}{8}$ or, equivalently, that $\operatorname{surp}(G)\geq\frac{\sqrt{8m+1}-1}{8}$, which is sharp when $G$ is a clique on an odd number of vertices.
@@
In general, if $G$ is a disjoint union of constantly many cliques, then its maximum cut is $m/2+O(\sqrt{m})$. This naturally raises the question: can one improve this bound if $G$ is far from being a disjoint union of cliques? One way to ensure that a graph is far from a disjoint union of cliques is to forbid some fixed graph $H$ as a subgraph. The study of the MaxCut, and in particular the surplus, in $H$-free graphs was initiated by Erdős and Lovász in the 1970’s [40]. A landmark result in the area is due to Alon [2], who proved that if a graph $G$ has $m$ edges and contains no triangles, then $\operatorname{surp}(G)=\Omega(m^{4/5})$, and this bound is tight. Two natural generalizations of this result are to consider graphs without short cycles, and graphs that avoid complete graphs $K_{r}$.
@@
The surplus in graphs without short cycles has been studied extensively [3, 6, 12, 46], with tight bounds obtained in [12, 46]. On the other hand, determining the minimum surplus in $K_{r}$-free graphs is much more difficult. Alon, Bollobás, Krivelevich, and Sudakov [3] proved that for every $r$, there exists $\varepsilon_{r}>0$ such that every $K_{r}$-free graph has surplus at least $m^{1/2+\varepsilon_{r}}$. This was improved by Carlson, Kolla, Li, Mani, Sudakov, and Trevisan [21], and further strengthened by Glock, Janzer, and Sudakov [46], who proved $\operatorname{surp}(G)=\Omega_{r}\big(m^{\frac{1}{2}+\frac{3}{4r-2}}\big)$. Nevertheless, these bounds seem far from optimal: Alon, Bollobás, Krivelevich and Sudakov conjectured in [3] that the true bound should be $\operatorname{surp}(G)\geq m^{3/4+\varepsilon_{r}}$ for some $\varepsilon_{r}>0$. This conjecture remains open. For many years, it was a tantalizing open problem to find any absolute constant $\varepsilon>0$, independent of $r$, such that every $K_{r}$-free graph has surplus $\Omega_{r}(m^{1/2+\varepsilon})$. Glock, Janzer and Sudakov [46] write “Arguably, the main open problem is to decide whether there exists a positive absolute constant $\varepsilon$ such that any $K_{r}$-free graph with $m$ edges has surplus $\Omega_{r}(m^{1/2+\varepsilon})$.” Our next main result not only resolves this problem, but also shows that one can guarantee such a surplus by forbidding extremely large cliques as well.
@@
There is a close relationship between the MaxCut of a graph $G$ and its smallest eigenvalue. It is well known that $\operatorname{surp}(G)\leq|\lambda_{n}|n$ (see e.g. Claim 5.1 for a short proof). A good way to think about the surplus is as a robust version of the smallest eigenvalue: in many natural cases $\operatorname{surp}(G)=\Theta(|\lambda_{n}|n)$, but $\operatorname{surp}(G)$ is much less sensitive to local modifications and harder to study.
The proof of Theorem 1.3 is presented in Section 10. In Appendix˜B, we discuss a variant of Theorem 1.3 which shows that graphs that are far from the disjoint union of cliques also have large surplus.
1.3 Smallest eigenvalue
A central topic of spectral graph theory is understanding the structure of graphs whose adjacency matrix has large (i.e. not very negative) smallest eigenvalue. Let $G$ be an $n$-vertex graph and let $\lambda_{n}$ denote the smallest eigenvalue of its adjacency matrix. Probably the best-known theorem in spectral graph theory involving the smallest eigenvalue is the celebrated Hoffman bound (see e.g. [50]), which states that $\lambda_{n}$ controls the independence number of the graph. In particular, if $G$ is an $n$-vertex $d$-regular graph, then $\alpha(G)\leq\frac{n|\lambda_{n}|}{|\lambda_{n}|+d}$. The Expander Mixing Lemma is of equal importance, stating that the maximum of $|\lambda_{n}|$ and the second-largest eigenvalue $\lambda_{2}$ determines the expansion and mixing properties of the graph [4]. Moreover, as we mentioned before, $\lambda_{n}$ controls the maximum cut.
A simple consequence of the Cauchy’s interlacing theorem is that if $G$ is non-empty, then $\lambda_{n}\leq-1$ with equality if and only if $G$ is the disjoint union of cliques. In the 1970’s, Cameron, Goethals, Seidel, and Shult [20] gave a complete characterization of graphs satisfying $|\lambda_{n}|\leq 2$, which are exactly generalized line graphs and some sporadic examples with at most 36 vertices. Koolen, Yang, and Yang [66] obtained a partial characterization in the case $|\lambda_{n}|\leq 3$ using integral lattices.
Beyond these specific values, much less is known. Hoffman [54] studied structural properties of graphs with $\lambda_{n}\geq-\lambda$, for some fixed constant $\lambda$, and his work was extended by Kim, Koolen, and Yang [60], who proved the following structure theorem for regular graphs satisfying $|\lambda_{n}|\leq\lambda$: one can find dense induced subgraphs $Q_{1},\dots,Q_{c}$ in $G$ such that each vertex lies in at most $\lambda$ of $Q_{1},\dots,Q_{c}$, and almost all edges are covered by the union of $Q_{1},\dots,Q_{c}$. However, the proof of this is based on certain forbidden subgraph characterizations and Ramsey-theoretic arguments, and the results are no longer meaningful if $\lambda$ grows faster than polylogarithmically in $n$.
For highly structured graphs, such as strongly regular graphs (SRGs), it is known [72] that if $|\lambda_{n}|$ is at most a small polynomial of the average degree, then the graph belongs to one of two special families (see also [65]). However, these results rely on the highly structured nature of SRGs. We refer the interested reader to the survey of Koolen, Cao, and Yang [64] for a general overview of the topic.
Many of these results show that the property of having small $|\lambda_{n}|$ and the existence of large trivial substructures, such as cliques, are interconnected. However, such results were previously only known when $|\lambda_{n}|$ is bounded by a constant, or growing very slowly with $n$. We prove that this phenomenon already starts to appear when $|\lambda_{n}|<n^{1/4-o(1)}$, and we show that graphs with smallest eigenvalue below this threshold are close to a trivial structure: the disjoint union of cliques. The exponent $1/4$ is also sharp, a celebrated construction of de Caen [30] related to equiangular lines provides a graph with smallest eigenvalue $|\lambda_{n}|=\Theta(n^{1/4})$ which is far from the disjoint union of cliques. We say that an $n$-vertex graph $G$ is $\delta$-close to some family of graphs $\mathcal{F}$ if one can change at most $\delta n^{2}$ edges to non-edges and vice versa to transform $G$ into a member of $\mathcal{F}$.